OpenStudy (anonymous):

how do you find the surface area to volume ratio? Surface Area: 1.5 Volume: 0.125

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

A ratio of two numbers is their quotient when the first number is divided by the second number (not zero).

OpenStudy (anonymous):

erh well i looked up on the internet and im quite confused. I was thinking you just do 1.5:0.123

OpenStudy (anonymous):

Could you show me how?

OpenStudy (skullpatrol):

$$\Huge \frac{1.5}{0.125}=\frac{\frac{3}{2}}{\frac{1}{8}}$$

OpenStudy (anonymous):

so... that would be the answer?

OpenStudy (anonymous):

what would it be as a decimal

OpenStudy (skullpatrol):

simplify the above expression

OpenStudy (anonymous):

3:4?

OpenStudy (skullpatrol):

The answer is 12

OpenStudy (anonymous):

because 3 multiplied by 4?

OpenStudy (skullpatrol):

yes

OpenStudy (anonymous):

super awesome! thank you. I gave you a medal.. need anything else?

OpenStudy (skullpatrol):

thanks for asking :)

OpenStudy (skullpatrol):